### v2

# Guardians and Offenders: A Survey on Harmful Content Generation and Safety Mitigation of LLM

Abstract.
Large Language Models (LLMs) have revolutionized content creation across digital platforms, offering unprecedented capabilities in natural language generation and understanding. These models enable beneficial applications such as content generation, question and answering (Q&A), programming, and code reasoning. Meanwhile, they also pose serious risks by inadvertently or intentionally producing toxic, offensive, or biased content. This dual role of LLMs, both as powerful tools for solving real-world problems and as potential sources of harmful language, presents a pressing sociotechnical challenge. In this survey, we systematically review recent studies spanning unintentional toxicity, adversarial jailbreaking attacks, and content moderation techniques. We propose a unified taxonomy of LLM-related harms and defenses, analyze emerging multimodal and LLM-assisted jailbreak strategies, and assess mitigation efforts, including reinforcement learning with human feedback (RLHF), prompt engineering, and safety alignment. Our synthesis highlights the evolving landscape of LLM safety, identifies limitations in current evaluation methodologies, and outlines future research directions to guide the development of robust and ethically aligned language technologies.
Large Language Models (LLMs), Harmful Language, Text Generation, Text Classification, Content Moderation, Prevention, Jailbreak
1. Introduction
The digital age has brought unprecedented connectivity, but with it, a surge in harmful online interactions has emerged. As we navigate this complex landscape, LLMs have emerged as a pivotal technology, uniquely positioned to both combat and paradoxically contribute to this challenge. On the positive side, LLMs offer a powerful tool for content moderation in the face of proliferating toxic speech, online harassment, offensive language, and biased narratives. Their remarkable capabilities of language understanding have been widely used to identify and flag problematic text. This can help overcome the limitations of traditional methods that often struggle with the complex understanding of online content.
However, these powerful text generation capabilities also make LLMs a potential source of harm, whether unintentionally (28, ; 29, ; 501, ; 534, ; 540, ) or deliberately (Chen2024-lv, ; Wang2024-zh, ; Mangaokar2024-ii, ; Jiang2023-ee, ; Liu2024-rd, ; Li2024-pj, ; Lin2024-ly, ; Gong2023-xr, ; Shayegani2023-in, ). LLMs can produce polarized comments or biased opinions in these circumstances, contributing to the very problem they could help solve if not carefully managed or aligned. This generation of LLM-created content can pose further risks, ranging from individual psychological distress (501, ; 534, ; 540, ) and exposure to misinformation to broader societal harms, such as the facilitation of malicious online campaigns or fraud (Gong2023-xr, ; Shayegani2023-in, ).
In this context, this literature review synthesizes recent research, focusing on the potential and challenges of LLMs in harmful text. To deepen our investigation, we formulate three core research questions (RQ1–RQ3).
  • RQ1. What is the current landscape of related studies?
  • RQ2. What are the challenges that LLMs bring, from unsafe performance to promising safety?
  • RQ3. What are the potential uses of LLMs in combating harmful generation and content?
The first research question surveys the current landscape of research on LLM safety, including model types, harm categories, and evaluation methodologies. The second research question addresses the multifaceted challenges in controlling harmful outputs, ranging from the unintended generation of toxic or biased content to sophisticated adversarial attacks, such as jailbreaks and multimodal exploits. The third explores the emerging capabilities of LLMs themselves to act as agents of safety through classification, moderation, self-monitoring, and preventive generation.
Our findings highlight a dual trajectory: On the one hand, LLMs remain susceptible to generating harmful content. However, LLMs also demonstrate significant potential to enhance safety by acting as classifiers, content moderators, counter-speech generators, and prevention systems capable of introspection and adaptation. This duality necessitates a shift in perspective: from viewing LLM safety as a static constraint to designing dynamic, context-aware, and self-correcting systems. Our synthesis offers a comprehensive foundation for future research in building reliable and resilient LLMs in high-stakes environments.
We organize this survey as follows: Section 2 introduces the role of LLMs about harmful content, including key definitions, terminology, and our data collection methodology. Section 3 provides a landscape analysis of existing research, summarizing the range of LLM models studied and the types of harmful content addressed. Section 4 examines the core challenges of harmful content generation by LLMs, including both unintentional outputs and those deliberately induced by malicious actors. Section 5 focuses on the potential of LLMs in detecting, moderating, and preventing harmful content. Finally, Section 6 presents our key findings, discusses broader implications, and outlines future research directions.
2. Background
2.1. Categorization of harmful content in LLM
We systematically categorize the harmful content generated by LLM into four types according to their impact mechanisms, as shown in Table 1. Toxic content encompasses both explicit abuse—such as profanity or slurs—and implicit hostility, which relies on rhetorical devices (e.g., euphemism, sarcasm, metaphor) and shared commonsense or social‐norm knowledge to convey harm  (284, ). Harassment refers to directed interpersonal attacks—bullying, doxxing threats, or group‐based verbal aggression—aimed at intimidating or silencing individuals or collectives (vidgen2019much, ). Offensive language spans explicit hateful remarks that are overt and relatively easy to detect, as well as implicit offenses whose harm emerges only in specific contexts or against particular targets (das2024offensivelang, ). Finally, bias content targets individuals or groups based on protected or socially salient attributes—such as race, gender, religion, ethnicity, or sexual orientation—and perpetuates systemic inequalities (xu2024hatespeech, ). Recognizing these categories is crucial for developing targeted detection and mitigation strategies within LLM.
Table 1. Definitions of content safety concerns across different categories.
Category Author and Year Definition
Toxic content Wen et al. (2023) (284, ) Explicit toxic content involves overtly abusive language such as swearwords. Implicit toxic content is conveyed through a variety of linguistic features (such as euphemism, sarcasm, circumlocution, metaphor) and extralinguistic knowledge (such as commonsense knowledge, world knowledge, and social norms).
Online Harassment Vidgen et al.(2019) (vidgen2019much, ) Interpersonal attacks, such as harassment and bullying, and verbal attacks against groups.
Offensive Language Das et al.(2024) (das2024offensivelang, ) Offensive language typically falls into two categories: explicit and implicit. Explicit offensive language is overtly hateful and generally easier to identify. In contrast, implicit offensive language is more subtle; while it may not directly convey hatred, it can still be perceived as offensive depending on the context and the sensitivity of the targeted group.
Biased Content Albladi et al.(2025) (xu2024hatespeech, ) Bias-based hate speech refers to harmful language that targets individuals or groups based on protected or socially salient identity attributes, such as race, gender, religion, ethnicity, or sexual orientation.
2.2. Terms and definitions
To ground our discussion of LLM safety, we introduce four central concepts—text generation, jailbreak attacks, content moderation, and text classification—that together span the lifecycle of interaction with harmful content. These definitions, drawn from both academic and industry sources (Table 2), provide a shared foundation for analyzing both the risks and safeguards associated with LLM deployment.
Text generation—known as natural language generation—involves producing coherent, human-like text by predicting and decoding sequences of tokens; effective decoding strategies elevate LLMs from mere next-token predictors to versatile generative engines (becker2024text, ; li2024pre, ; LLM-text-generation, ). Text classification is the task of assigning predefined labels—such as sentiment or topic—to an input text, forming the backbone of many downstream filtering and analysis pipelines  (sun2023text, ; dong2022survey, ). Content moderation refers to the systematic review and filtering of user-generated content to enforce a platform’s policies and community standards, preventing abuse and facilitating cooperation  (tspa2025content, ; grimmelmann2015virtues, ). Finally, Jailbreak attacks exploit carefully crafted prompts to bypass an LLM’s built-in safety guardrails, coaxing prohibited or harmful outputs and posing a serious security challenge for both research and commercial deployments (Zhou2024-uh, ; Yu2024-pm, ; Chen2023-eq, ). This layered framework enables a comprehensive understanding and systematic addressing of LLM safety challenges, from the fundamental generation mechanisms to the practical implementation of safety measures, while acknowledging both defensive capabilities and potential vulnerabilities. In this paper, we define Prevention as the tool or process to prevent the generation of the toxic content, online harassment, offensive language and biased content, which are defined in Table 1.
Table 2. Definitions of LLM tasks related to harmful content generation and mitigation.
Category Author and Year Definition Origin
Text Generation Becker et al. (2024) (becker2024text, ) Text generation, also known as natural language generation, is a computational task focused on producing coherent human-like text from structured or unstructured non-linguistic inputs. Academia
Li et al. (2022) (li2024pre, ) The central aim of text generation is to develop an end-to-end framework that learns to map input representations to fluent textual outputs, minimizing the need for manual rule design. Academia
Nabi et al. (2024) (LLM-text-generation, ) Language models are trained to anticipate subsequent tokens in a sequence, and the strategy used to choose among possible continuations—known as decoding—plays a key role in transforming them from predictive engines into task-effective text generators. Industry
Text Classification Wang et al.(2023) (sun2023text, ) Text classification is a classical NLP task that assigns a category label to a given piece of text, and is commonly used for sentiment analysis, topic detection, and more. Academia
Min et al.(2023)(dong2022survey, ) Text classification is the task of assigning pre-defined labels (e.g., sentiment, topic) to a given text. Academia
Content Moderation Trust & Safety Professional Association (2023) (tspa2025content, ) Content moderation involves actively reviewing user-generated content to ensure it aligns with a platform’s guidelines. Industry
Grimmelmann. (2015) (grimmelmann2015virtues, ) Moderation functions as a mechanism to promote cooperative engagement and to safeguard against misuse and abusive conduct. Academia
Jailbreak Zhou et al.(2024) (Zhou2024-uh, )
Yu et al.(2024) (Yu2024-pm, )
Jailbreak attacks involve inducing LLMs to generate harmful content by circumventing safety guardrails through specially crafted prompts, posing a significant threat to LLM security by eliciting content that was originally designed to be prohibited. Academia
Chen et al.(2023) (Chen2023-eq, ) Jailbreaking refers to the phenomenon where carefully crafted prompts elicit harmful responses from models, bypassing their built-in content security measures and safety alignment, which persists as a significant challenge for commercial LLMs. Industry
2.3. Data Preparation
OpenAlex (priem2022openalex, ), an open-source repository of scholarly metadata, is used to collect related studies. It has two advantages. First, it allows us to gather both recently archived preprints and published articles from journals and conferences. Second, OpenAlex offers its free and open querying dataset, ensuring our analysis can be replicated within the community without licensing barriers. Table 3 shows the paper collection and filtering keywords. The search terms include two major categories, including LLMs, and terms implying harmful content, as shown in Table 3.
The initial search with the OpenAlex yields 3347 records up to $May.22^{nd},2025$. Before the primary screening, 218 duplicate records are removed, along with an additional 260 records identified as potential duplicates. This preprocessing step results in 3087 records for the screening phase. During this stage, a significant portion, specifically 2236 records, is excluded as they are determined to be irrelevant to the scope of this survey. We use GPT-4 as a judge to support this relevance check. Following screening, 751 reports are sought for full-text retrieval and eligibility assessment. Of these, 379 reports are subsequently excluded. The primary exclusion criteria applied at this stage are: (1) the study is not related to the unsafe aspects of LLMs, or (2) the study lacks technical details and elaborated evaluation.
After this multi-stage filtering process, 372 studies were deemed eligible and included in the final review. Figure 1 presents a visual distribution of these studies across three dimensions: technical approach, LLM model type, and hate speech category. Subsequently, we conducted a manual review for each of the 372 papers to extract key information for further analysis.
Table 3. Keywords of Filtering the Papers
Category Terms
LLMs GPT, ChatGPT, GPT-4, GPT-3, Gemini, Claude, PaLM, Palm 2, Llama, Llama 2, Mistral, Mixtral, Qwen, Falcon, davinci, BLOOM, LaMDA, Titan, Anthropic, Open AI, Meta AI, Google AI, DeepMind, Bard, Large language model, Large Language Models, LLM, decoder-only models
Harms profanity, profane language, racism, racial bias, racial discrimination, sexism, gender bias, gender discrimination, sexual harassment, homophobia, online hate, toxic comment, hate speech, cyberbullying, toxicity, abusive, trolling, harassment, content moderation, toxic, offensive, hateful, discrimination, harmful content, extremist content, harmful language, radicalization, online abuse, online harassment, extremist
Figure 1. Distribution of Related Studies by Technical Approach, Model Type, and Hate Speech Category. From left to right, the three polar bar plots display the number of research papers categorized by: (1) technical approach; (2) model type; and (3) type of hate speech. Each bar’s height indicates the number of papers in that category. Colors represent the primary application domain: Content Moderation, Evaluation, Jailbreak, Prevention, Text Classification, or Text Generation. The inner rings summarize the total number of studies highlighting the ”Good” or ”Bad” aspects of LLMs in the context of hate speech. For example, papers under Text Generation may explore both beneficial and harmful impacts of LLMs related to hate speech.
3. RQ1. What is the landscape of current studies?
This section presents an overview of the collected studies, highlighting trends in research focus, methodologies, and key findings.
Figure 1 presents a multidimensional overview of the existing literature on LLMs in the context of harmful content, organizing the collected research papers along three axes: technical approach, model type, and harmful content types. This comprehensive landscape reveals both the offensive behaviors of LLMs, in which they act as generators of harm, and their evolving capabilities as guardians against unsafe content. Each polar plot includes inner rings denoting whether a study primarily examines the harmful or mitigative role of LLM, offering a high-level view of the dual emphasis of the field.
The first polar chart categorizes the studies by the technical approaches. Here, prompt engineering and fine-tuning dominate the methodological space, particularly in tasks related to text generation, content moderation, and classification. These methods offer accessible, post-hoc means of shaping model behavior without requiring architectural changes or retraining from scratch. In contrast, fewer studies explore adversarial learning or architectural redesign, reflecting a community preference for efficient, low-friction interventions. In particular, prompt-based strategies are frequently employed in both offensive and defensive contexts, underscoring their flexibility in eliciting or suppressing harmful outputs.
The second chart breaks down the research by model type. GPT-based models account for the largest share of studies, especially in jailbreak detection and safety evaluations, likely due to their commercial prominence and wide accessibility. LLaMA (touvron2023llama, ), Claude, and open-source transformer models such as RoBERTa (liu2019roberta, ) and BERT (devlin2019bert, ) also feature prominently in classification and moderation tasks. However, many region-specific or emerging models (e.g., Gemini (team2023gemini, ; team2024gemini, ), Baichuan (li2024baichuan, ), Qwen (bai2022training, )) remain underexplored, revealing a geographical and institutional bias in LLM safety research. This skew suggests a concentration of risk analysis and mitigation strategies in a narrow model family, raising concerns about generalization and global robustness.
The third chart classifies studies by the type of harm addressed: identity-based hate, offensive language, toxic behavior, and online harassment. Most work focuses on lexical or semantic toxicity, such as slurs, stereotypes, and profanity, which are easier to label and benchmark. Identity-based harms (e.g., racial or gender bias) are also widely studied, often through fairness evaluations. However, complex and relational harms, such as online harassment, remain less explored, despite their prevalence in the real world. This gap highlights the need for broader harm taxonomies and more context-aware moderation frameworks.
Crucially, each dimension in Figure 1 reveals a notable asymmetry: while harmful capabilities of LLMs are thoroughly documented, a greater portion of the literature explores their potential as safety enablers. Across technical methods, model types, and harm categories, studies emphasizing the guardian role of LLMs outnumber those treating them as offenders (e.g., 287 ”Good” vs. 163 ”Bad” in technical methods, 294 vs. 162 in model types, and 340 vs. 209 in harm types). These findings align with the dual framing of this survey. Section 5 elaborates on the persistent risks posed by LLMs, including unintentional generation of toxic content, embedded bias, and vulnerability to adversarial exploits. And Section 6 examines how the very same models can be repurposed as self-correcting, context-sensitive agents for safety.
4. RQ2. What are the challenges that LLMs bring to the harmful generation?
The challenges of LLMs in addressing hate speech manifest in both unintentional generations and intentional exploits, presenting a complex landscape of security concerns.
Unintentional Generation of Harmful Content
Despite significant advancements in safety alignment, including mitigation techniques such as RLHF (bai2022training, ) and prompt filtering (kumar2023certifying, ), LLMs remain susceptible to generating hate speech. Research has shown that even well-aligned models like GPT-4 (28, ; 29, ; 501, ; 534, ; 540, ), Gemini (598, ), and LLaMA (29, ) can still produce toxic, biased, or offensive content under certain conditions. This inherent vulnerability highlights the fundamental challenge in achieving robust safety guarantees in LLMs.
Intentional Exploitation and Attack Evolution
Beyond unintentional generation, a more concerning challenge lies in the intentional exploitation of LLMs through jailbreak attacks—strategically engineered prompts that circumvent safety mechanisms to induce restricted responses, such as the generation of hate speech. These attacks have demonstrated a clear evolution pattern: text-based adversarial methods (Chen2024-lv, ; Wang2024-zh, ; Mangaokar2024-ii, ), compositional and stealthy approaches (Jiang2023-ee, ; Liu2024-rd, ; Li2024-pj, ; Lin2024-ly, ), and multi-modal strategies (Gong2023-xr, ; Shayegani2023-in, ) that exploit vulnerabilities in cross-modal processing pipelines. Most alarmingly, LLMs themselves can now be leveraged to generate adversarial prompts, effectively enabling self-jailbreaking or assisting in jailbreak attacks on other models. Techniques like the Tree of Attacks with Pruning (TAP) demonstrate how attacker LLMs can autonomously refine jailbreak prompts for maximum toxicity and evasion. These evolving techniques reveal critical weaknesses in LLM training, decoding, and alignment mechanisms—enabling the generation of increasingly sophisticated harmful content.
Table 4. Evolution of LLM Challenges in Hate Speech Generation: A systematic categorization of challenges in controlling harmful content, from unintentional vulnerabilities to evolving attack strategies, supported by representative research findings.
Category Key Findings & Characteristics Works
Unintentional Generation
Safety Alignment Limitations Despite RLHF and prompt filtering, models exhibit inherent vulnerabilities in content control (bai2022training, ), (kumar2023certifying, )
Model Behavior Well-aligned models (GPT-4, Gemini, LLaMA) produce toxic content under certain conditions (28, ; 29, ), (598, )
Bias Manifestation Systematic biases in generated content across gender, race, and cultural dimensions (501, ; 534, ; 540, )
Intentional Exploitation
Basic Text Attacks Adversarial techniques focusing on prompt manipulation (Chen2024-lv, ; Wang2024-zh, )
Advanced Composition Sophisticated methods combining multiple attack vectors and stealthy approaches (Jiang2023-ee, ; Liu2024-rd, )
Multi-modal Integration Cross-modal attacks exploiting visual-language processing:
- Typographic visual attacks (FigStep)
- Image-text combinations (Jailbreak in Pieces)
- Role-play based attacks (VRP)
(Gong2023-xr, ), (Shayegani2023-in, )
LLM-Assisted Generation Self-evolving attack capabilities:
- Tree of Attacks with Pruning (TAP)
- In-context learning for jailbreak
- Iterative refinement
(268, ; Deng2023-vd, )
Summary of the Challenge Landscape
As summarized in Table 4, these challenges present a two-fold threat to LLM safety: (1) the unintentional generation of hate speech (Section 4.1) (including safety alignment limitations, model behavior issues, and bias manifestation), (2) the advancement of jailbreak strategies for intentional abuse (evolving from basic text attacks to advanced multi-modal integration), and (3) the emergent ability of LLMs to generate jailbreak prompts (Section 4.2) through self-evolving capabilities. This combination of inherent vulnerabilities and evolving attack capabilities, as illustrated by the trend in Table 4, underscores the significant and multifaceted challenges LLMs face in addressing hate speech.
4.1. Unintentional Hate Speech Generation by LLMs
Despite continuous advancements in ethical alignment, research consistently shows that LLMs still generate hate speech. Studies analyzing models such as GPT-4, Llama, and Claude have identified LLMs can produce toxic content (3, ; 16, ; 19, ; 28, ; 29, ), offensive language (91, ; 501, ; 534, ; 540, ; 605, ), and biases related to gender (103, ; 598, ; levy-etal-2024-gender, ) and race (91, ; 501, ; 605, ). Notably, these occurrences are often unintentional, emerging as a byproduct of the model’s learning process (501, ; 540, ; 605, ; 610, ). This issue arises possibly because LLMs learn linguistic associations from their training corpora, which contain such harmful content that inadvertently reinforces the generation of harmful outputs (16, ; 29, ; 91, ).
Figure 2. Unintentional and intentional hate speech generation by LLMs
Several studies demonstrate how LLMs can unintentionally generate toxic content. A recent study shows that LLMs trained on massive, web-sourced datasets, which often contain harmful content, can perpetuate toxic behaviors (mendu2025safer, ). While GPT-4 exhibits reduced toxicity compared to earlier models, it can still produce toxic responses (28, ). Similarly, Wang et al. (2025) reveal that RLHF-trained models with ”personification,” although enhancing user interactivity, continue to generate toxic outputs. This indicates that current safety measures remain insufficient (wang2025explor, ). Furthermore, studies on LLM applications in text generation for transgender and non-binary (TGNB) individuals show that models frequently misgender them and can embed toxic language in responses to gender disclosures (16, ).
LLMs have also been found to generate offensive language. A study on AI-generated news content indicates that LLMs can amplify offensive stereotypes, particularly in their representation of marginalized communities (501, ). Similarly, an evaluation of ChatGPT’s moderation framework shows that while it attempts to filter offensive language, its censorship mechanisms sometimes overcompensate, which results in selective filtering that inadvertently suppresses certain perspectives (540, ). Another study explores how LLMs handle humor generation. It reveals that LLMs often struggle with culturally offensive languages involving minorities (534, ). In cybersecurity applications, LLMs have been shown to generate offensive code payloads, enabling the creation of malicious tools that facilitate cyber attacks (91, ).
Bias remains one of the most persistent and concerning issues in LLMs, such as gender and racial bias. A study on LLM-generated letters of recommendation finds that models consistently favor male-coded language (103, ). In scientific writing, LLMs replicate and sometimes amplify gendered stylistic differences, with male-authored abstracts being more directive and receiving higher citations, whereas female-authored abstracts emphasize insight and emotion, often undervalued in academic discourse (598, ). Moreover, research on implicit biases in LLMs demonstrates how models reinforce gender and racial stereotypes through word associations and narrative structures, when they often favor dominant demographic groups (605, ). A study on relationship decision-making finds that LLMs systematically favor female-associated perspectives, which challenges traditional male dominance but simultaneously introduces new biases (levy-etal-2024-gender, ).
4.2. Intentional Strategies for Hate Speech Generation
Jailbreaking mechanisms in LLMs exploit model vulnerabilities to bypass safety measures and generate harmful outputs. These attacks have demonstrated a clear evolutionary trajectory: from simple text-based manipulations to compositional strategies, and ultimately to sophisticated multimodal approaches, as illustrated in Figure 3. This progression not only reflects the the growing sophistication of adversarial threats but also parallels the rapid advancement of modern LLMs, particularly as they integrate more advanced features like multimodal understanding and cross-modal reasoning.
Figure 3. Intentional strategies for hate speech generation. a. Adversarial suffix embeddings (Wang et al. (2024) (Wang2024-zh, )); b. Multi-turn dialogue (Chen et al. (2023) (34, )); c. Disguise malicious content (Liu et al. (2024) (Liu2024-rd, )); d. Sensitive word substitution (Ren et al. (2024) (742, )); e. Multi-modal attack (Shayegani et al. (2023) (Shayegani2023-in, )); f. LLM-assisted attack (Deng et al. (2023) (Deng2023-vd, )).
4.2.1. Text-based Adversarial Attack Strategies:
Initial methods focused on manipulating textual inputs to induce unsafe outputs from LLMs with minimal modifications. Chen et al. (2024) proposed RL-JACK (Chen2024-lv, ), a reinforcement learning-enabled black-box attack that optimizes prompt generation and refinement strategies. Wang et al. (2024) introduced ASETF (Wang2024-zh, ), leveraging gradient-based optimization of adversarial suffix embeddings to achieve high attack success rates with reduced computational cost. Mangaokar et al. (2024) developed PRP (Mangaokar2024-ii, ), which introduces universal adversarial prefixes to bypass guard mechanisms across various LLM architectures. These early-stage strategies laid the foundation for more advanced and structured attack methodologies.
Another effective approach is multi-turn adversarial prompting, where an attacker gradually escalates a conversation to induce toxic behavior. Chen et al. (2023) revealed that single-turn filtering mechanisms often fail in multi-turn exchanges, where initial non-toxic queries evolve into offensive outputs as chatbots adapt to the conversation flow (34, ). They proposed a ToxicChat method which demonstrate that 82% of toxic chatbot response can be induced by seemingly benign multi-turn interactions, bypassing moderation due to the contextual shifts. Zhou et al. (2024) showed that LLMs may fail to reject cautionary or borderline unsafe queries, particularly within multi-turn conversations where each turn subtly contributes to an overarching malicious intent (zhou2024speak, ). Fan et al. (2024) proposed that multi-turn dialogues can lead to more severe outcomes in hate speech generation, driven by increased conversational complexity and the compounding effects of bias accumulation (Fan2024-Fa, ).
4.2.2. Compositional and Stealthy Attack Strategies:
More sophisticated techniques embed harmful intentions within seemingly benign prompts. Jiang et al. (2023) presented Prompt Packer (Jiang2023-ee, ), a method that utilizes compositional instruction attacks by embedding malicious content within complex writing or dialogue tasks. Liu et al. (2024) proposed DRA (Liu2024-rd, ), which disguises harmful instructions and guides LLMs to reconstruct the malicious content in their outputs. Li et al. (2024) introduced DrAttack (Li2024-pj, ), a framework that decomposes harmful prompts into multiple sub-prompts and recombines them through in-context learning. Lin et al. (2024) developed ABJ (Lin2024-ly, ), leveraging LLMs’ reasoning capabilities to embed harmful queries under analytical tasks. These methods highlight the shift toward semantic camouflage and hidden intent transmission. In addition, Deshpande et al. (2023) and Giorgi et al. (2025) proposed a persona-assignment strategy. By instructing LLMs to adopt specific identities, these attacks cause models to reinterpret otherwise harmful prompts as benign within the context of the assigned persona. This technique significantly amplifies toxicity, with some personas increasing harmful outputs by up to sixfold (66, ; giorgi2025humanllmbiaseshate, ).
Recent work has demonstrated that sensitive word substitution can effectively conceal harmful intent and induce LLMs to generate hate speech while evading detection. Bianchi et al. (2024) proposed a bait-and-switch technique, which manipulates safe LLM-generated outputs post-generation through automated word substitutions or adversarial transformations. This approach enables adversaries to produce harmful content at scale while maintaining plausible deniability, as the original text remains superficially compliant with moderation filters (369, ). Similarly, Ren et al. (2024) introduced Feign Agent Attacks (F2A), which bypass content moderation by embedding fake security validation signals within prompts and obfuscating toxic content in Python code—splitting offensive text into variables or disguising it through string operations. This allows the LLM to reconstruct and output harmful content during execution while circumventing detection mechanisms (742, ).
4.2.3. Multi-modal Attack and LLM-Assisted Attack Strategies:
Recent research has demonstrated increasingly sophisticated attack strategies, evolving from traditional text-based approaches to complex multimodal attacks and even LLM-assisted generations. These strategies can be categorized into two main branches:
Multi-modal and Cross-Modal Attacks. Recent research extends attacks beyond text inputs, integrating multimodal elements such as images and visual cues. Gong et al. (2023) proposed FigStep (Gong2023-xr, ), which targets Vision-Language Models (LVLMs) through typographic visual attacks that exploit weaknesses in cross-modal processing. Shayegani et al. (2023) introduced Jailbreak in Pieces (Shayegani2023-in, ), combining image and text inputs to reveal vulnerabilities in multimodal LLMs (MLLMs). Liu et al. (2025) presented PiCo to introduce a gray-box attack method that generates adversarial images or audio to convey specific harmful instructions to MLLMs (Liu2025-Pi, ). Building on these approaches, Ma et al. (2024) introduced the innovative ”Role-play” concept through Visual Role-play (VRP)—a method that uses high-risk character images and benign instructions to mislead MLLMs into generating malicious responses, achieving significantly higher attack success rates than prior approaches (ma2024-vi, ).
LLM-Assisted Attack Generation. In a concerning development, LLMs themselves have become tools for generating sophisticated attack prompts. This represents a significant escalation in attack capabilities, as LLMs can now systematically assist in bypassing content moderation in other models (268, ; Deng2023-vd, ). The Tree of Attacks with Pruning (TAP) methodology exemplifies this approach, using an attacker LLM to iteratively refine jailbreak prompts while an evaluator LLM assesses their effectiveness. This sophisticated technique has demonstrated over 80% success rates against advanced models like GPT-4 and Gemini-Pro, even in the presence of robust moderation tools (268, ). Further advancing this field, researchers have explored in-context learning for jailbreak prompt generation, training LLMs to mimic and refine adversarial input patterns (Deng2023-vd, ). Starting from a small set of manually crafted prompts, these systems can develop broader, more effective attack sets through iterative refinement, often surpassing human-created prompts in their success rates.
Table 5. Evolution of LLM Jailbreak Strategies
Strategy Type Key Techniques Representative Methods
Text-based Reinforcement learning optimization RL-JACK (Chen2024-lv, )
Gradient-based suffix embedding ASETF (Wang2024-zh, )
Universal adversarial prefixes PRP (Mangaokar2024-ii, )
Compositional Complex task embedding Prompt Packer (Jiang2023-ee, )
Disguised instruction reconstruction DRA (Liu2024-rd, )
Sub-prompt decomposition DrAttack (Li2024-pj, )
Analytical task embedding ABJ (Lin2024-ly, )
Multimodal Typographic visual attacks FigStep (Gong2023-xr, )
Combined image-text inputs Jailbreak in Pieces (Shayegani2023-in, )
5. RQ3. What are the potential uses of LLMs in combating harmful generation and content?
With LLMs’ advanced language understanding, they are increasingly recognized for their potential. We organize this section based on the tasks they perform. First, they are used to build datasets (das2024offensivelang, ; 392, ; 383, ) or generate counter-speech (hong-etal-2024-outcome, ; 105, ; 263, ; 422, ). Second, LLMs’ reasoning capability can help classify diverse forms of harmful content (masud2023focal, ; nasir2023llms, ; dutta2023modeling, ; lee2024mitigating, ; xu2024hatespeech, ; Agarwal2023-ml, ; ovalle2023you, ) and perform content moderation (lee2024exploring, ; wang2023evaluating, ; he2024you, ; vishwamitra2024moderating, ; lu2024toxic, ; wang2023self, ; ovalle2023you, ; zhang2024genderalign, ). Finally, previous research efforts show that LLMs can contribute to the prevention  (xie2023empirical, ; ji2023beavertails, ; lee2023square, ; bang2023enabling, ; qureshi2024refine, ; kaneko2024evaluating, ; wang2024model, ; jorgensen2023improving, ) and the defense against jailbreaking threats (Phute2023-ue, ; Chen2023-eq, ; kumar2023certifying, ; beck25sensitivity, ).
Then, we explore the potential of LLMs in two areas: addressing existing harmful content in the first three subsections, and preventing the possible harmful generation in the latter two subsections, respectively. Cross several key tasks, namely mitigating existing harmful content through generation (Section 5.1), identifying and classifying harmful content (Section 5.2), content moderation (Section 5.3), and prevention (Section 5.4)—while simultaneously examining crucial strategies for the defense against jailbreak and the prevention of LLM-related risks (Section 5.5).
5.1. Mitigating Existing Harmful Content through Generation
Researchers have leveraged LLMs to generate synthetic hate speech datasets, which enable the development of more effective detection/classification models and reduce reliance on ethically sensitive real-world data. LLMs also contribute to counter-speech generation, where LLM responses effectively combat online hate speech in real-time. Additionally, LLMs can be used for red-teaming other LLMs, where controlled hate speech generation helps stress-test moderation systems and expose vulnerabilities in content filtering mechanisms. The potentials of LLMs’ text generation in hate speech are thus demonstrated from these three aspects.
5.1.1. Hate Speech Generation for Dataset Construction
LLMs have been increasingly used to generate synthetic hate speech datasets, providing valuable resources for hate speech detection, classifier evaluation, and safety alignment. These datasets differ in language coverage, dataset construction methods, and target applications, ensuring a comprehensive and diverse approach to detecting and mitigating toxic content.
Several studies focus on generating explicit and implicit hate speech datasets using LLMs. OffensiveLang contains 8,270 ChatGPT-generated offensive and non-offensive texts spanning 38 target groups across race, gender, religion, and disability, designed to enhance classifier robustness against implicit toxicity (das2024offensivelang, ). Similarly, K-OMG produces culturally aligned offensive language datasets for Korean. It leverages prompt-based LLM generation to address low-resource language detection challenges (392, ). ParaFusion generates 3.5 million paraphrases and filters offensive content to improve lexical and syntactic diversity in hate speech detection to ensure a balanced dataset for NLP applications (383, ).
Datasets targeting bias and stereotypes are also constructed using LLMs. SeeGULL Multilingual introduces 25,861 LLM-generated stereotypes across 20 languages and provides a benchmark for evaluating fairness and bias mitigation in hate speech detection (397, ). Meanwhile, a GPT-generated dataset annotated for gender bias quantifies bias degrees in offensive text, offering fine-grained insights into identity-based attacks and biased language patterns (286, ).
LLM-generated datasets cover various forms of hate speech, from explicit slurs to coded language and microaggressions. Some datasets focus on identity-based hate speech, targeting racial, gender, and religious biases (das2024offensivelang, ), while others address sarcasm and indirect hate speech. KoCoSa, a Korean Context-aware Sarcasm Detection Dataset, integrates LLM-generated sarcastic dialogues to train hate speech detection models on contextual toxicity (kim-etal-2024-kocosa, ). Certain datasets focus on multilingual and cultural variations in hate speech. IndicLLMSuite provides LLM-generated toxic prompts across 22 Indic languages, enabling multilingual hate speech detection (394, ). Similarly, SeeGULL Multilingual captures geo-culturally situated stereotypes, which offers insights into hate speech variations across different regions (397, ).
Hate speech datasets generated by LLMs are widely applied in benchmarking detection models and stress-testing moderation systems. GPT-HateCheck generates functional hate speech test cases. These test cases address the limitations of template-based benchmarks like HateCheck and ensure more nuanced and real-world hate speech samples (jin-etal-2024-gpt-hatecheck, ).
Counterfactually Augmented Data (CAD) expands hate speech datasets with adversarial variations, improving classifier generalization across different hate speech contexts (283, ). CAD-based dataset augmentation has been explored for security-related text classification. It has shown potential in hate speech detection, spam detection, and fraud detection (223, ).
LLM-generated datasets also play a role in safety alignment and ethical model training. KoTox, a Korean Toxic Instruction Dataset, contains 39,000 toxic instruction-output pairs to help LLMs reject harmful queries, strengthening ethical AI alignment (264, ). Additionally, IndicLLMSuite generates toxic prompts and non-toxic responses for Indic languages, contributing to hate speech detection efforts in underrepresented linguistic contexts (394, ).
5.1.2. LLMs for Counter Speech Generation
Counter speech has emerged as a key strategy for mitigating online hate speech, and recent advances in LLMs have opened new possibilities for automating its generation. Researchers have explored various approaches to improve the effectiveness of LLM-generated counter-speech. They focus on aspects such as conversation outcome constraints (hong-etal-2024-outcome, ), discourse structuring (105, ), instructional prompting (263, ), and zero-shot performance analysis (422, ).
Hong et al. (2024) propose to enhance LLM-generated counter speech by optimizing for conversation outcomes. It uses reinforcement learning and instruction prompting to encourage civil engagement and reduce incivility in social media discussions (hong-etal-2024-outcome, ). Another approach proposed by Hedderich et al. (2024) applies LLM-powered chatbots for cyberbullying intervention. Thus, teachers are allowed to design interactive training tools where students practice counter-speech in a safe, role-playing environment (105, ).
Beyond these applications, a discourse-aware framework (DisCGen) improves counter-speech generation by incorporating Segmented Discourse Representation Theory (SDRT), which leads to more diverse and contextually relevant responses (263, ). A zero-shot evaluation comparing GPT-2, DialoGPT, ChatGPT, and FlanT5 further demonstrates that ChatGPT produces the most effective counter-speech, but larger models risk generating more toxic responses. The study by Saha et al. (2024) also finds that manual prompting outperforms other strategies, though LLMs struggle with humor and rhetorical counter-speech (422, ).
5.1.3. Hate Speech Generation for Red-Teaming
Red-teaming LLMs plays a critical role in identifying vulnerabilities and evaluating safety mechanisms, particularly in the context of hate speech generation. By systematically probing models with adversarial prompts, researchers can stress-test moderation systems and enhance content safety mechanisms. Recent studies have explored a variety of automated red-teaming techniques leveraging LLM text generation, ranging from reinforcement learning-based attacks to optimization-driven adversarial auditing.
One approach is reinforcement learning-based adversarial prompt generation, as demonstrated by the three-stage red-teaming framework (Exploration, Establishment, and Exploitation), which systematically manipulates LLMs to increase their rate of toxic output, exposing hidden vulnerabilities that standard safety techniques fail to address (27, ). Another strategy is discrete optimization-based auditing, where frameworks like ARCA use log-probability approximations to search for high-risk input-output pairs, making them more effective than manual red-teaming methods (123, ).
More recent methods leverage LLM-based adversarial auditing. TroubleLLM introduces text style transfer and adversarial conditioning techniques. It generates highly effective adversarial prompts that outperform human-crafted and template-based red-teaming methods, which makes it an automated solution for systematically testing model vulnerabilities (385, ). Similarly, GFlowNet fine-tuning with MLE smoothing enables the creation of diverse and transferable attack prompts. The generated adversarial testing is more robust and adaptable compared to traditional reinforcement learning-based strategies (520, ).
5.2. Identifying and Classifying Harmful Content
Table 6. Summary of LLM Text Classification
Strategy Type Model Key Techniques References
Training-Based BERT, RoBERTa, LLaMA Supervised adaptation on hate speech datasets; domain-specific optimization (masud2023focal, ; nasir2023llms, ; dutta2023modeling, )
GPT-3.5, GPT-4 RLHF for ethical alignment (lee2024mitigating, )
LLaMA, Mistral Robust fine-tuning with adversarial examples to resist prompt attacks (zhou2024multimodal, ; wu2024exploring, )
Prompt-Based GPT-3, LLaMA Zero-shot / few-shot learning using task-specific prompt templates (xu2024hatespeech, ; Agarwal2023-ml, ; ovalle2023you, )
GPT-4 Output filtering or reformulation of toxic generations (e.g., via contrastive decoding or classifiers) (Zhang2024, ; cao2023toxic, )
Bias Mitigation BERT, LLaMA Counterfactual data augmentation, debiasing objectives, demographic analysis (You2024, ; Liu2024, ; sun2024bias, )
Data-Centric GPT-3.5, GPT-4 LLM-generated hate speech and counter-speech data augmentation (lee2024toxicity, ; meguellati2025llm, )
Cross-Lingual & Context mBERT, XLM-R, LLaMA Multi-language model adaptation for low-resource hate speech classification (alenezi2024detecting, )
RoBERTa, LLaMA Coreference resolution and contextual history modeling for implicit hate (perez2023bias, ; xu2024hatespeech, )
Evaluation-Based All models Perplexity-based scoring, calibration curves, benchmark comparison (jansen2022perplexed, ; li2024detecting, ; lim2023evaluating, )
Despite significant progress in text recognition capabilities of LLM in recent years, hate speech remains a complex and evolving challenge in cyberspace. Currently, the main method for detecting such harmful content is text classification, which assigns structured labels such as ”harmful”, ”harmless”, ”hateful”, ”insulting”, etc. to unstructured text. Owing to their strong contextual reasoning and ability to generalize from examples, LLMs have demonstrated exceptional performance in this task. Nevertheless, these models also introduce new concerns, including the amplification of pre-existing social biases, difficulty in detecting subtle or implicit forms of hate, and vulnerability to adversarial manipulation.
This section presents a structured examination of text classification methods for hate speech detection. We begin by examining how fine-tuning techniques enhance model adaptability and detection accuracy. It then discusses the role of diverse model architectures in improving contextual understanding and multilingual detection capabilities. In addition, we explore prompt engineering strategies to reduce biases and annotation costs while improving interpretability. To address dataset scarcity and class imbalance, we discuss data augmentation and synthetic text synthesis. While cross-lingual transfer learning leverages multilingual transformers to extend detection to low-resource languages. Subsequently, we analyze the contributions of diverse model architectures to enhance contextual sensitivity and enable robust performance in multilingual settings.
5.2.1. Characterizing Toxicity
Defining and measuring toxicity present inherent challenges. Luo et al. (luo2023legally, ) highlighted a potential misalignment between LLM-generated hate speech and established legal definitions, pointing to inconsistencies with regulatory goals. Furthermore, the perception and manifestation of harmful content can vary culturally. Lee et al. (lee2024exploring, ) demonstrated this by collecting datasets showing variability in toxicity norms across different English-speaking regions, including North America, Australia, Singapore, and the United Kingdom. Evaluating the trustworthiness of LLMs requires a multi-faceted approach. Qian et al. (qian2024towards, ) proposed evaluating trustworthiness across five dimensions—reliability, toxicity, privacy, fairness, and robustness—using linear probing. Their findings suggest that a model’s capacity for stable and secure generation can often be discerned even in early pretraining checkpoints.
5.2.2. Prompt Engineering
Prompt engineering has become instrumental for enhancing model transparency, interpretability, and fairness in hate speech detection. Carefully crafted prompts explicitly guide model reasoning, significantly reducing annotation efforts and improving accuracy and ethical compliance by explicitly instructing models to avoid stereotypical biases. For example, structured prompt templates have been shown to reduce the generation of harmful stereotypes (dorner2022human, ; rosenblatt2022critical, ). Identity aware prompting mitigates fairness-conditioned prompts improve safety compliance (Agarwal2023-ml, ; ovalle2023you, ). Context-sensitive prompts have been shown to significantly enhance transparency and alignment with human values (Som2023-yq, ). Moreover, advanced prompting methods, including zero shot and few shot learning, further facilitate generalization to novel hate speech patterns and implicitly toxic content, enhancing model robustness and reducing potential biases in real-world deployments (xu2024hatespeech, ; Zhang2024, ).
5.2.3. Fine tuning and Alignment for Hate Speech Detection
Fine-tuning LLMs remains one of the most effective strategies for domain adaptation in hate speech detection. By adjusting pretrained models to task-specific objectives and sensitive content distributions, fine-tuning enhances recognition of implicit toxicity, subtle insults, and identity directed hate. For instance, label-sensitive fine-tuning explicitly exposes underrepresented hate cases, improving performance on rare but critical examples (tuck2024unmasking, ), while staged pipelines progressively capture nuanced forms of aggression across diverse datasets (kumar2024towards, ). To address class imbalance, a focal loss strategy reweights examples based on classification difficulty, yielding substantial gains on toxic content benchmarks (masud2023focal, ), and curriculum learning further boosts robustness by gradually exposing the model to increasingly challenging toxic inputs (nasir2023llms, ).
In multi-label settings, frameworks that leverage inter-label correlations and hierarchical dependencies enable finer-grained detection of overlapping toxic behaviors (dutta2023modeling, ). Complementary studies have evaluated how pretraining choices and dataset shifts influence fine-tuning outcomes (rossi2023investigating, ), and benchmarks of noise robustness and distributional perturbations have been established (garcia2023benchmark, ). To mitigate fairness concerns and domain drift, demographic-sensitive training and episodic fine-tuning tailor models to underrepresented groups and evolving content distributions (gupta2023toxicity, ; hickey2025peripatetic, ), while experiments on Arabic toxic corpora highlight ongoing challenges in multilingual generalization (almohaimeed2023thos, ). Generative replay mechanisms have been introduced during fine-tuning to retain ethical constraints and prevent catastrophic forgetting (pendzel2024generative, ). Finally, hybrid approaches that combine fine-tuning with prompt engineering, such as prompt-based supervision to resist adversarial jailbreaks (zhou2024multimodal, ) and dynamic prompting to guide models through ambiguous or subversive inputs (Gupta2024, ) alongside reconfiguration techniques for identifying harmful intent and suppressing undesirable outputs (wu2024exploring, ), further enhance both robustness and ethical alignment.
5.2.4. Data Augmentation and Text Synthesis for the Imbalance in Speech Categories
Data augmentation techniques and synthetic data generation address dataset scarcity and class imbalance in hate speech detection. Back-translation and paraphrasing via pretrained models can produce diverse rephrasings of toxic and non-toxic content, effectively expanding training distributions (kolla2022study, ). Prompt-based synthesis with large generative models has been shown to generate realistic hate and non-hate examples, improving robustness to novel linguistic patterns (maity-etal-2024-toxvidlm, ). Embedding-based matching further refines augmented samples to maintain semantic coherence, boosting generalization across varied expressions of toxicity (gupta2024toxicity, ).
5.2.5. Cross-lingual Transferability for Classification
Cross-lingual transfer learning extends detection models to resource-scarce languages by leveraging shared multilingual representations. Multilingual transformers fine-tuned on high-resource datasets can be adapted to Arabic, Hindi, and other low-resource contexts with minimal labeled data, yielding significant performance gains (alenezi2024detecting, ). Coreference-driven approaches explicitly model referential chains (pronouns, named entities) within text, capturing subtle, identity-targeted hate that relies on context and anaphora. Integrating coreference resolution modules into transformer backbones has been demonstrated to improve detection accuracy in nuanced scenarios (perez2023bias, ).
5.2.6. Evaluation and Benchmarking for Robustness and Fairness
Rigorous evaluation and benchmarking methodologies ensure reliable, fair, and transparent performance in hate speech detection models. Benchmarking studies systematically evaluate models across diverse hate speech datasets, explicitly addressing issues of bias, fairness, and generalization capabilities (vogel2023explaining, ). Comprehensive surveys guide the selection and development of robust metrics and datasets, fostering improved model reliability and fairness across contexts (lim2023evaluating, ). Furthermore, specialized benchmarking datasets explicitly tailored for hate speech enhance robustness evaluations, accommodating diverse linguistic variations and nuanced offensive language patterns (cao2023toxic, ). Calibration approaches refine confidence estimations, crucially enhancing model trustworthiness in sensitive applications (li2024detecting, ). Joint dataset creation and benchmarking initiatives facilitate consistent model comparisons and accelerate knowledge transfer, improving hate speech detection performance across research communities (hossain2024misgendermender, ). Additionally, established safety taxonomies categorize prompt-related risks systematically, further supporting the safe and ethical deployment of models (hu2023automated, ).
5.3. Content Moderation
The rapid proliferation of LLMs, particularly those built upon Transformer-based architectures, has completely changed the application of natural language processing. However, with these advances, there are growing concerns that LLMs may generate harmful content, either inadvertently or through adversarial prompting (zhang2023joint, ). As a result, ensuring the safe and responsible deployment of LLMs has become a pressing research priority, with content moderation emerging as a critical area of focus. Recent advancements demonstrate that with careful prompt design, fine-tuning, or parameter editing, LLMs can perform moderation at scale while preserving language diversity and generation fluency. This subsection surveys technical approaches that transform LLMs into effective content moderators across monolingual and multilingual environments.
5.3.1. Prompt Engineering
Prompt engineering offers a means of controlling LLM outputs. Wang et al. (wang2023evaluating, ) used GPT-3 itself, prompted appropriately, to generate explanations for online content moderation decisions. He et al. (he2024you, ) explored detoxification in GPT-2 and T5 using prompt tuning. Vishwamitra et al. (vishwamitra2024moderating, ) developed a reasoning-based Chain-of-Thought (CoT) prompting technique that automatically generates and updates prompts for moderation purposes. A more direct approach involves vocabulary manipulation; Lu et al. (lu2024toxic, ) proposed pruning subword tokens associated with toxic words within trained LLMs. This simple mechanism demonstrated improved dialogue diversity and toxicity reduction in models like Llama-3.1_8B (lu2024toxic, ).
5.3.2. Alignment, and Parameter Editing
Researchers are actively developing methods to mitigate these risks, broadly categorized into safety training and safeguards, often employed concurrently (wang2023self, ). Several approaches focus on refining the models during or after training. Fine-tuning has proven effective in moderating specific biases, such as those related to pronouns or race (ovalle2023you, ). Alignment techniques aim to steer model behavior towards desired norms. Zhang et al. (zhang2024genderalign, ) generated a specialized alignment dataset to mitigate gender bias in LLMs specifically. Similarly, Direct Preference Optimization (DPO) trained on English preference data has shown promise in reducing toxicity even in multilingual generation across models like mGPT-1.3B, BLOOM, Llama3, and Aya-23 (li2024preference, ). Contrastive adversarial gender debiasing leverages contrastive learning techniques, reportedly outperforming standard adversarial debiasing methods (torres2024contrastive, ).
Beyond extensive retraining or fine-tuning, modifying specific model parameters offers a more targeted approach. Wang et al. (Wang2024-ac, ) discovered that editing only a small subset of model parameters can effectively modulate and detoxify generation while preserving the LLM’s general capabilities. Building on this, Elesedy et al. (elesedy2024lora, ) employed low-rank adapters (LoRA), demonstrating effective content moderation with significantly fewer (100x) guardrail parameters compared to full fine-tuning, preventing degradation of the generative model and suggesting potential for moderated models on resource-constrained devices like mobile phones (elesedy2024lora, ).
Unlearning techniques aim to selectively remove specific knowledge or behaviors, such as copyrighted or harmful content, from a trained model without degrading its overall utility. Negative Preference Optimization (NPO) is recognized as a state-of-the-art unlearning method. Fan et al. (fan2024simplicity, ) introduced a simplified preference optimization method for unlearning that removes the need for a reference model, potentially making the process more efficient (fan2024simplicity, ).
5.3.3. Cross-Lingual Considerations
Toxicity and its mitigation are not confined to a single language. Interestingly, safety alignment efforts can exhibit cross-lingual transfer. Li et al. (li2024preference, ) found that DPO training solely with English data could reduce toxicity in multilingual outputs. Lu et al. (lu2024every, ) further corroborated this multilingual potential, showing that unsafe generation patterns can influence outputs across languages. Crucially, their experiments indicated that unlearning methods applied by addressing harmful responses in both English and the content’s original language could effectively eliminate toxic generation across multiple languages supported by the model (lu2024every, ).
In summary, addressing toxicity in LLMs involves understanding its nuanced nature across contexts and languages and developing a diverse toolkit of mitigation strategies, ranging from comprehensive retraining and alignment to targeted parameter editing, sophisticated prompt engineering, and selective unlearning techniques.
5.4. Potential in Harm Prevention
Beyond filtering or correcting harmful content after generation, an emerging area of research focuses on preventing such content from arising in the first place. In this proactive role, LLMs are equipped with architectural safeguards, guided prompts, or ethical preference structures that steer their responses toward safer outputs instead of unintentional harm or deliberately crafted adversarial responses. These preventive strategies reflect a shift in focus from post hoc moderation to anticipatory control, minimizing the risk of offensive completions, biased responses, or policy-violating outputs before they surface. Through the integration of human preferences, bias-aware reasoning, and gated decoding mechanisms, large language models (LLMs) can be fine-tuned or guided to minimize the generation of unsafe or socially undesirable content. This subsection examines how safety can be embedded into model behavior by design, focusing on targeted interventions at the data, prompt, and parameter levels.
Xie et al. (xie2023empirical, ) employ counterfactual data augmentation (CDA) as a bias mitigation strategy for pretrained language models, specifically GPT-2 and BERT. Beavertails (ji2023beavertails, ) constructs a comprehensive Human-Preference Dataset comprising 333,963 question–answer pairs and 361,903 expert-labeled comparison instances, annotated for helpfulness and harmlessness. This dataset contributes significantly to enhancing LLM safety through RLHF. Lee et al. (lee2023square, ) builds a Korean language dataset of sensitive questions with corresponding acceptable and unacceptable responses. And the experiments show the dataset improves the acceptable responses for HyperCLOVA and GPT-3.
Some additional components are helpful to prevention. Bang et al. (bang2023enabling, ) introduce an additional fine-tuned classifier performing value-aligned supervision, suggesting that the value-based classifier improves inclusivity and explainability in LLM. REFINE-LM (qureshi2024refine, ) involves a simple model added on top of LLMs, which is cheaper to train with reinforcement learning and reduces biases while preserving the performance.
Prompt engineer contributes to preventing the generation of harmful content. Kaneko et al. (kaneko2024evaluating, ) uses Chain-of-Thought to evaluate the step-by-step generation of LLMs on bias, and the result shows that CoT prompting reduces this unconscious social bias in LLMs.
In efforts to improve LLM against adversarial attacks and malicious signals, Wang et al. (wang2024model, ) introduced model parameter editing to enhance robustness. Similarly, Jorgensen et al. (jorgensen2023improving, ) used activation-based methods, in which the mean activation of a target dataset is subtracted from the overall training activations, yielding effective steering vectors for the prevention of toxicity.
Another significant advancement in prevention techniques is Controllable Text Generation (CTG). Kim et al. (kim2023gta, ) proposed Gated Toxicity Avoidance (GTA), a CTG method that enables LLMs to steer clear of toxic topics, tones, and expressions. However, toxicity prevention remains a complex challenge, particularly when safety fine-tuning leads to unintended consequences. Zhao et al. (zhao2024learning, ) identified a phenomenon in which safety fine-tuning can inadvertently degrade model performance by causing the forgetting of downstream data. To counteract this issue, they introduced ForgetFilter, a pre-finetuning technique designed to preserve model capabilities while improving safety. (jain2023bring, ) tried self-supervised evaluation to monitor the performance of LLMs, which complements the labeled data provided by humans. Through supervised fine-tuning, SteelLM (dong2023steerlm, ) grants user control over model responses during inference, while delivering performance competitive with numerous RLHF-based state-of-the-art models. Besides, Antidote (huang2024antidote, ) aims to defend against the fine-tuning attacks. They detect and prune detrimental parameters that emerge as a result of the fine-tuning process, which prevents harmful generation while maintaining performance.
5.5. Defense against LLM Jailbreaking
The current landscape of LLM jailbreak research has witnessed the development of increasingly sophisticated attack and defense mechanisms. These approaches range from fundamental text-based manipulations to advanced multimodal strategies, reflecting the evolving complexity of adversarial threats against LLMs. Defense mechanisms have similarly evolved, incorporating adaptive self-monitoring, efficient safeguards, and context-aware strategies to enhance model security. This technical progression demonstrates both the increasing sophistication of attack methodologies and the corresponding advancement in protective measures, highlighting the ongoing challenge of maintaining LLM safety while preserving model utility.
5.5.1. Adaptive and Self-monitoring Defenses
Adaptive defense mechanisms leverage LLMs’ internal monitoring capabilities and systematic evaluation frameworks to provide real-time and certified protection. Phute et al. (2023) (Phute2023-ue, ) introduced the LLM Self Defense approach, enabling models to screen their own outputs for harmful content without additional fine-tuning. Chen et al. (2023) (Chen2023-eq, ) proposed Moving Target Defense strategies, employing randomized selection from multiple LLM instances to minimize attack success rates, reducing them from 37.5% to 0%. Furthermore, Kumar et al. (2023) (kumar2023certifying, ) developed the erase-and-check framework, offering certified safety guarantees against adversarial prompting with 92% detection accuracy. Wang et al. (2023) (Wang2023-vh, ) and Rao et al. (2023) (Rao2023-pr, ) provided comprehensive frameworks for systematic safety evaluation, using detailed taxonomies and standardized testing to assess potential vulnerabilities effectively.
5.5.2. Scalable Safeguards and Context-aware Defenses
Efficient defense mechanisms aim to minimize computational overhead while maintaining robust security. Kim et al. (2023) (Kim2023-mj, ) developed the Adversarial Prompt Shield (APS), a lightweight classifier using only 25% of the parameters compared to traditional solutions, achieving state-of-the-art performance. Additionally, these efficient defenses align with broader evaluation frameworks that prioritize both computational efficiency and model robustness, ensuring that security measures do not compromise processing speed or system scalability. Developing effective defenses requires addressing social and contextual factors that influence model behavior. Beck et al. (2023) (beck25sensitivity, ) investigated sociodemographic prompting, revealing variations in model responses based on demographic attributes, thereby underscoring the need for nuanced and socially informed safety mechanisms. These findings emphasize the importance of integrating social considerations into defense strategies, ensuring equitable and contextually appropriate model behavior across diverse user demographics.
5.5.3. Evaluation and Mechanism Analysis of Jailbreaking
Effective evaluation and interpretability are critical for understanding LLM vulnerabilities and developing robust defense strategies. In terms of safety evaluation, Wang et al. (2023) (Wang2023-vh, ) introduced the Do-Not-Answer dataset, offering a comprehensive benchmark for evaluating model responses to high-risk prompts across various scenarios. This dataset provides a hierarchical risk taxonomy covering multiple harm categories, enabling systematic assessments of LLM safety. Building on this foundation, Yu et al. (2024) (Yu2024-pm, ) developed a holistic framework that characterizes and evaluates LLM reliability across diverse jailbreak strategies, incorporating key metrics such as attack success rates, toxicity scores, and grammatical accuracy to provide a multi-dimensional view of LLM robustness. These evaluation frameworks have been further enhanced by recent works such as Chen et al. (2024) (Chen2024-lv, ), Wang et al. (2024) (Wang2024-zh, ), and Mangaokar et al. (2024) (Mangaokar2024-ii, ), which provide additional metrics and benchmarks for assessing different attack strategies. Complementing these evaluation frameworks, mechanism analysis has revealed deeper insights into LLM vulnerabilities. Zhou et al. (2024) (Zhou2024-uh, ) conducted a detailed analysis of hidden states in LLMs, demonstrating how jailbreak prompts circumvent safety mechanisms by disrupting intermediate representations rather than directly attacking initial ethical recognition. This understanding has been enriched by studies from Jiang et al. (2023) (Jiang2023-ee, ), Liu et al. (2024) (Liu2024-rd, ), Li et al. (2024) (Li2024-pj, ), and Lin et al. (2024) (Lin2024-ly, ), which explore various compositional attack mechanisms. Further theoretical work by Shen et al. (2024) (Shen2024-fe, ) in Mission Impossible proved that adversarial vulnerabilities are inevitable under current training paradigms, while research from Gong et al. (2023) (Gong2023-xr, ) and Shayegani et al. (2023) (Shayegani2023-in, ) extended these insights to multimodal contexts. The defensive aspects of these mechanisms have been explored by Phute et al. (2023) (Phute2023-ue, ), Chen et al. (2023) (Chen2023-eq, ), Kim et al. (2023) (Kim2023-mj, ), and Kumar et al. (2023) (kumar2023certifying, ), highlighting the importance of comprehensive assessment approaches that consider both empirical performance and theoretical limitations. These combined insights from both evaluation frameworks and mechanism analyses suggest the need for new training strategies to enhance robustness against adversarial prompting.
Table 7. Technical Classification
Technical Approach Methods Key Techniques Papers
Reinforcement Learning-based RL-JACK RL-based prompt optimization
Black-box attack framework
Dynamic refinement strategies
Chen et al.(2024)(Chen2024-lv, )
Wang et al.(2024)(Wang2024-zh, )
Kumar et al.(2023)(kumar2023certifying, )
Gradient Optimization ASETF
PGD
GradAttack
Adversarial suffix embeddings
Gradient-based optimization
Low computational cost
Token-level perturbation
Wang et al.(2024)(Wang2024-zh, )
Mangaokar et al.(2024)(Mangaokar2024-ii, )
Li et al.(2024)(Li2024-pj, )
Universal Adversarial PRP
UAP
AdvPrefix
Universal adversarial prefixes
Cross-architecture attack
Guard mechanism bypass
Transfer learning attack
Mangaokar et al.(2024)(Mangaokar2024-ii, )
Jiang et al.(2023)(Jiang2023-ee, )
Liu et al.(2024)(Liu2024-rd, )
Compositional Attacks Prompt Packer
DRA
DrAttack
ABJ
Complex task embedding
Instruction disguising
Sub-prompt decomposition
Task-based hiding
Semantic camouflage
Jiang et al.(2023)(Jiang2023-ee, )
Liu et al.(2024)(Liu2024-rd, )
Li et al.(2024)(Li2024-pj, )
Lin et al.(2024)(Lin2024-ly, )
Beck et al.(2023)(beck25sensitivity, )
Multimodal Attacks FigStep
Jailbreak in Pieces
VisualJail
Typographic attacks
Image-text fusion
Cross-modal vulnerabilities
Visual prompt injection
Gong et al.(2023)(Gong2023-xr, )
Shayegani et al.(2023)(Shayegani2023-in, )
Wang et al.(2023)(Wang2023-vh, )
Yu et al.(2024)(Yu2024-pm, )
Defense Mechanisms LLM Self Defense
Moving Target Defense
APS
Erase-and-check
SafeGuard
Self-monitoring screening
Random instance selection
Lightweight classification
Safety certification
Adaptive protection
Real-time detection
Phute et al.(2023)(Phute2023-ue, )
Chen et al.(2023)(Chen2023-eq, )
Kim et al.(2023)(Kim2023-mj, )
Kumar et al.(2023)(kumar2023certifying, )
Wang et al.(2023)(Wang2023-vh, )
Zhou et al.(2024)(Zhou2024-uh, )
Shen et al.(2024)(Shen2024-fe, )
Recent technical approaches in LLM jailbreaking have demonstrated increasing sophistication and diversity. Reinforcement learning-based methods, represented by RL-JACK (Chen2024-lv, ), optimize attack strategies through trial-and-error interactions with target LLMs, using reward signals based on attack success rates and output toxicity levels. This black-box approach proves particularly effective against commercial LLMs where model parameters are inaccessible. Gradient-based methods like ASETF (Wang2024-zh, ) focus on optimizing adversarial inputs through gradient computations, achieving efficient attacks by directly manipulating the embedding space while maintaining low computational requirements. Building upon these foundations, universal adversarial methods such as PRP (Mangaokar2024-ii, ) develop attack patterns that work across multiple LLM architectures, creating universal adversarial prefixes that can be prepended to various inputs to induce harmful outputs. More sophisticated compositional attacks have emerged, breaking down harmful prompts into seemingly benign components. For instance, Prompt Packer (Jiang2023-ee, ) embeds malicious content within complex writing tasks, DRA (Liu2024-rd, ) employs disguised instructions, DrAttack (Li2024-pj, ) splits harmful prompts into multiple innocent-looking sub-prompts, and ABJ (Lin2024-ly, ) conceals attack components within analytical tasks. The frontier has further expanded to multimodal attacks, with FigStep (Gong2023-xr, ) exploiting typographic elements in images and Jailbreak in Pieces (Shayegani2023-in, ) combining image and text inputs to bypass traditional safety measures. In response to these evolving threats, defense mechanisms have been developed, including LLM Self Defense (Phute2023-ue, ) for continuous self-monitoring, Moving Target Defense (Chen2023-eq, ) for dynamic model selection, APS (Kim2023-mj, ) for lightweight classification of potential attacks, and the Erase-and-check framework (kumar2023certifying, ) for certified safety guarantees.
6. Discussion
Our comprehensive analysis of LLM challenges in hate speech generation reveals a complex and rapidly evolving landscape that demands immediate attention from multiple stakeholders. The findings demonstrate a clear progression in both the nature and sophistication of challenges: from unintentional vulnerabilities inherent in model training to increasingly sophisticated intentional exploits, culminating in the emergence of self-evolving attack capabilities. These challenges manifest across multiple dimensions, from technical limitations in safety alignment to complex social implications of model biases. The rapid evolution of attack methodologies, particularly in multi-modal integration (Gong2023-xr, ; Shayegani2023-in, ) and LLM-assisted generation (268, ; Deng2023-vd, ), suggests that current safety measures may be insufficient for future threats. Our analysis yields three critical areas for discussion: (1) key findings regarding the technical and social dimensions of these challenges, (2) implications for stakeholders across the technical, industry, and policy domains, and (3) future research directions needed to address emerging threats. In the following sections, we examine each of these areas in detail, highlighting the urgent need for more robust and adaptive approaches to ensuring LLM safety in social media contexts.
6.1. Key Findings
Our comprehensive review reveals several critical findings about the challenges and evolution of LLM safety in social media contexts, highlighting the increasing complexity and urgency of addressing these issues:
Dual Nature of Challenges. As demonstrated in Table 4, the challenges in controlling harmful content generation exhibit a dual nature that demands a comprehensive understanding and response. The unintentional vulnerabilities persist despite significant advances in safety alignment techniques like RLHF (bai2022training, ) and prompt filtering (kumar2023certifying, ). Even the most sophisticated and well-aligned models like GPT-4, Gemini, and LLaMA demonstrate concerning patterns of systematic biases across gender, race, and cultural dimensions (501, ; 534, ; 540, ). Such biases are exhibited in diverse ways, from subtle linguistic preferences to more overt stereotypical associations, affecting everything from professional content generation to casual interactions (91, ; 501, ; 605, ).
The intentional exploitation landscape presents an even more alarming picture, with attack strategies evolving at an unprecedented pace. The progression from basic text-based attacks (Chen2024-lv, ; Wang2024-zh, ) to sophisticated multi-modal integration methods (Gong2023-xr, ; Shayegani2023-in, ) demonstrates not only the growing complexity of security challenges but also the increasing creativity of adversarial approaches. Recent studies have shown that these advanced attack methods can achieve success rates significantly higher than traditional approaches, with some techniques increasing harmful outputs by up to sixfold (66, ; giorgi2025humanllmbiaseshate, ).
Accelerating Technical Evolution. The technical sophistication of attacks has shown remarkable acceleration, following a clear evolutionary pattern that challenges existing safety paradigms. Initial text-based adversarial techniques focused on simple prompt manipulation (Chen2024-lv, ; Wang2024-zh, ), but quickly evolved to more advanced compositional methods (Jiang2023-ee, ; Liu2024-rd, ) that exploit subtle weaknesses in model architectures. The emergence of multi-modal attacks represents a significant leap in attack complexity, with methods like FigStep’s typographic visual attacks (Gong2023-xr, ) and VRP’s role-play based approaches (ma2024-vi, ) demonstrating the vulnerability of cross-modal processing pipelines.
Most concerningly, the development of LLM-assisted generation methods has opened a new frontier in attack sophistication. The Tree of Attacks with Pruning (TAP) method (268, ) has achieved unprecedented success rates of 80% against advanced models like GPT-4 and Gemini-Pro, even in the presence of robust moderation tools. This self-improving attack capability is particularly alarming as it demonstrates how LLMs can be weaponized against their own safety measures. Furthermore, research has shown that these automated attack methods often surpass human-created prompts in their effectiveness (Deng2023-vd, ), suggesting a fundamental challenge to current safety paradigms.
Interconnected Impact Dimensions. The challenges manifest across multiple interconnected dimensions, creating a complex web of technical, social, and security implications. In the technical dimension, fundamental limitations in model alignment persist despite significant advances in safety techniques (bai2022training, ; kumar2023certifying, ). These limitations are not merely engineering challenges but reflect deeper questions about the nature of language model training and control. Studies have shown that even models trained with explicit safety constraints can generate harmful content when presented with carefully crafted inputs (31, ; 34, ; 66, ; 94, ; 108, ; 284, ; 369, ).
The social dimension reveals deep-rooted issues with bias and fairness that go beyond simple content filtering. Research has demonstrated systematic biases in generated content (501, ; 534, ; 540, ), with particular concerns around gender representation (103, ; 598, ; levy-etal-2024-gender, ), racial bias (91, ; 501, ; 605, ), and cultural stereotypes. These biases often emerge in subtle ways, making them particularly challenging to detect and mitigate. Studies have shown that these biases can be amplified in specific contexts, such as professional writing or news generation (91, ; 501, ).
The security dimension demonstrates an evolving arms race between attack methods and defense mechanisms. Recent research has shown that attackers are increasingly leveraging the models’ own capabilities for more effective exploits (268, ; Deng2023-vd, ). This includes sophisticated techniques for prompt manipulation, context poisoning, and even the generation of adversarial examples that can bypass traditional safety measures. The emergence of multi-turn dialogue exploits (34, ; 744, ) and semantic conditioning techniques (94, ; 108, ) suggests that attack methods are becoming increasingly sophisticated and harder to defend against.
6.2. Implications
The findings from our study have significant implications across multiple stakeholder groups, requiring coordinated responses at various levels:
Technical Implications. Model development must undergo fundamental transformations to address the identified challenges. Current safety alignment techniques, while valuable, have shown clear limitations (bai2022training, ; kumar2023certifying, ). Future architectures need to incorporate more robust safety mechanisms that can address both unintentional biases and intentional attacks. This includes developing new approaches to model training that can better handle the complexity of human values and social contexts.
The success of multi-modal attacks (Gong2023-xr, ; Shayegani2023-in, ) necessitates the development of integrated defense systems that can protect against cross-modal vulnerabilities. This requires not only technical innovations in model architecture but also new approaches to data curation and training. Evaluation frameworks must expand beyond traditional metrics to assess implicit biases and potential for harmful content generation (501, ; 534, ). Recent research suggests the need for more sophisticated testing methodologies that can capture subtle forms of bias and potential for harm (das2024offensivelang, ; jin-etal-2024-gpt-hatecheck, ).
Furthermore, the emergence of LLM-assisted attacks (268, ; Deng2023-vd, ) calls for new paradigms in model security. Traditional approaches to content filtering and safety boundaries may no longer be sufficient when models can actively participate in their own exploitation. This suggests the need for more dynamic and adaptive security measures that can evolve alongside potential threats.
Industry Implications. Companies deploying LLMs must fundamentally rethink their approach to safety and security. The demonstrated vulnerabilities in current systems (28, ; 29, ) require more than incremental improvements to existing safety measures. Organizations need to implement comprehensive safety protocols that include continuous monitoring systems capable of detecting both traditional and emerging attack patterns.
The rise of sophisticated attacks like TAP (268, ) and multi-modal exploitation methods (Gong2023-xr, ; ma2024-vi, ) necessitates the development of adaptive defense mechanisms that can evolve alongside threat patterns. This includes investing in research and development of new safety technologies, as well as establishing clear protocols for incident response and mitigation.
Industry stakeholders must also address the ethical implications of their technology. Studies showing systematic biases (501, ; 534, ; 540, ) suggest the need for more rigorous testing and validation procedures before deployment. Companies should establish clear guidelines for responsible AI deployment, incorporating lessons learned from documented exploitation cases and ensuring transparency in their safety measures.
Policy Implications. The regulatory framework must evolve to address the complex challenges presented by LLMs in social media contexts. Studies showing systematic biases (501, ; 534, ; 540, ) call for stronger oversight of model training and deployment practices. This includes developing standards for model evaluation and certification, as well as establishing clear accountability mechanisms for when systems fail to meet safety requirements.
The rapid evolution of attack methods (268, ; Deng2023-vd, ) suggests the need for flexible regulatory approaches that can keep pace with technological advancement. Policy makers should focus on developing comprehensive ethical guidelines that address both current vulnerabilities and emerging threats. This includes establishing frameworks for regular safety audits and impact assessments, as well as requirements for transparency in model deployment and operation.
The demonstrated potential for societal harm through both unintentional bias (501, ; 534, ) and intentional exploitation (Gong2023-xr, ; Shayegani2023-in, ) underscores the importance of proactive policy measures. This includes developing guidelines for responsible AI development and deployment, establishing mechanisms for public oversight and accountability, and ensuring that regulatory frameworks can adapt to emerging challenges. Recent research on privacy preservation and user protection suggests the need for comprehensive policies that protect individual rights while enabling beneficial AI applications.
6.3. Future work
As the LLM landscape evolves rapidly, future efforts should address both the limitations of current safety alignment techniques and the untapped potential of LLMs in handling hate speech. We organize the future work into two major thrusts: (1) overcoming the challenges of LLMs in hate speech, and (2) amplifying the potentials of LLMs in this domain.
6.3.1. Overcoming the Challenges of LLMs in Hate Speech
Leveraging LLM-Generated Hate Speech for Beneficial Applications: A counterintuitive yet promising direction is the controlled use of LLM-generated hate speech for constructive purposes. Although filtering toxic content is necessary, overly restrictive moderation may hinder research and the development of effective defenses. Recent “green teaming” studies suggest that selectively bypassing filters in secure settings can yield valuable societal applications (140, ).
One such application is generating synthetic hate speech datasets, which help train more robust detection models without exposing annotators to harmful real-world content. Additionally, red-teaming exercises that use controlled hate speech to probe moderation systems can uncover hidden vulnerabilities, enabling stronger alignment safeguards.
Moreover, LLMs show promise in generating counter speech—automated, context-aware responses that challenge hateful language. Future research ought to concentrate on creating adaptive counter speech models that respond dynamically to evolving hate speech patterns while adhering to ethical standards.
Investigating the Underlying Mechanisms of Hate Speech Generation: To develop more effective alignment techniques, it is essential to understand how and why LLMs generate hate speech. Studies indicate that current safety methods may only suppress surface-level behavior without modifying the internal representations responsible for toxicity (liu-etal-2024-intrinsic, ).
Future research should investigate the effects of alignment strategies—both intrinsic (e.g., during training) and extrinsic (e.g., through RLHF)—on the model’s internal token dynamics and knowledge representations. Techniques like model probing and adversarial testing can shed light on whether harmful associations are truly suppressed or merely masked. In parallel, ethical reasoning and value alignment should be embedded earlier in the model development pipeline, rather than relying solely on post-hoc interventions (691, ; 246, ).
Preventing and Defending Against Jailbreak Attacks: Future work should focus on developing more robust defense frameworks that can resist advanced attack strategies. Mangaokar et al. (2024) (Mangaokar2024-ii, ) emphasize the need for improving guard model architectures to better handle universal adversarial prefixes. Kumar et al. (2023) (kumar2023certifying, ) recommend enhancing safety alignment as a potential alternative to RLHF. With the rise of multimodal LLMs, enhancing cross-modal safety alignment is crucial. Gong et al. (2023) (Gong2023-xr, ) stress the importance of developing defenses for Vision-Language Models (LVLMs), while Shayegani et al. (2023) (Shayegani2023-in, ) suggest improving alignment techniques to mitigate vulnerabilities across diverse input modalities.
Advanced detection mechanisms are essential for identifying jailbreak attempts in real-time. Phute et al. (2023) (Phute2023-ue, ) propose integrating in-context learning and logit biasing to reinforce safety responses in LLM outputs. Kim et al. (2023) (Kim2023-mj, ) advocate for refining adversarial prompt datasets and improving detection robustness through better dataset curation. Improving LLMs’ understanding of user intent is a crucial defense strategy. Jiang et al. (2023) (Jiang2023-ee, ) recommend enhancing intent recognition and command disassembly mechanisms to counter compositional instruction attacks. Li et al. (2024) (Li2024-pj, ) suggest extending prompt decomposition techniques to multilingual contexts and integrating them with other adversarial frameworks for better attack prevention.
Balancing Performance and Safety in LLMs: A significant challenge in LLM development lies in striking an effective balance content safety and model performance. While stricter safety mechanisms reduce the risk of generating harmful content, they often lead to diminished linguistic creativity and response diversity (548, ). For instance, studies on RLHF have shown that heavily aligned models frequently exhibit ”mode collapse”, which produces overly deterministic outputs that reduce syntactic and semantic variation. This constraint is particularly problematic in applications that demand creative or context-sensitive generation, such as dialogue systems or narrative writing.
Further compounding this issue, Bianchi et al. (2023) (bianchi2023safety, ) observed that excessive fine-tuning for safety can result in over-cautious behavior, where models unjustifiably refuse benign prompts merely because they resemble unsafe ones. Similarly, Ma et al. (2023) (ma2023adapting, ) found that extensive supervised fine-tuning and data engineering can cause LLMs to overreact to innocuous inputs. These findings underscore a critical usability-safety tension, where over-correction hampers legitimate functionality and undermines user trust.
To address this, future research should explore controlled text generation techniques that allow for nuanced moderation without severely restricting the expressive capabilities of LLMs. One promising direction is the use of dynamic attribute graphs (DATG), which provide a structured way to modulate the occurrence of key words while preserving generative diversity (liang-etal-2024-controlled, ). Such approaches can enable fine-grained safety controls that adapt to contextual nuances rather than enforcing blanket content prohibitions.
6.3.2. Amplifying the Potentials of LLMs in Hate Speech
Enhancing the Capability in Hate Speech Classification: Despite notable advances in using LLMs for hate speech detection and classification, several areas demand further exploration to ensure scalable and equitable deployment.
First, reducing reliance on large labeled datasets remains a priority. Future work should investigate few-shot, zero-shot, and parameter-efficient learning techniques—such as adapters, LoRA, and prompt tuning—that maintain performance under class imbalance and data scarcity (xu2024hatespeech, ; nasir2023llms, ).
Second, hate speech often spans multiple modalities and domains. Integrating multimodal fusion and domain-adaptation techniques can help models generalize across text, image, video, and platform-specific contexts (alenezi2024detecting, ; Cam2023-nn, ). Likewise, achieving multilingual fairness requires combining cross-lingual transfer with fairness-aware training to avoid performance disparities in underrepresented languages and dialects (xu2024hatespeech, ).
Third, robustness against adversarial perturbations, such as manual jailbreaks and automated prompt attacks, must be improved. Hybrid defenses combining prompt engineering, adversarial training, and continual learning will be crucial to maintaining classifier integrity over time (wu2024exploring, ; zhou2024multimodal, ).
Finally, to promote real-world adoption and trust, future systems should incorporate interpretability tools, calibrated uncertainty estimates, and transparent moderation pipelines (vogel2023explaining, ; li2024detecting, ). Standardized benchmarks that reflect diverse linguistic, cultural, and demographic contexts are equally important for ensuring reproducibility and ethical progress (lim2023evaluating, ; cao2023toxic, ).
Enhancing the Capability in Hate Speech Prevention: Preventing the generation of hate speech requires not only robust detection/classification systems but also proactive strategies that guide LLMs towards safer and more inclusive outputs. A growing body of research emphasizes the importance of enriched supervision to strengthen safety alignment and bias mitigation through larger, more diverse, and value-aligned pre-trained datasets (xie2023empirical, ; ji2023beavertails, ; lee2023square, ; bang2023enabling, ; qureshi2024refine, ).
One fundamental challenge that limits the capability of LLMs in hate speech prevention is the narrow scope of human preferences represented in existing labeled data. Ji et al. (2023) (ji2023beavertails, ) observe that demographic homogeneity among annotators can embed a limited worldview into model training, which undermines generalization and fairness. Likewise, Lee et al. (2023) (lee2023square, ) and Dong et al. (dong2023steerlm, ) argue that value assumptions embedded in certain languages may constrain LLMs’ normative boundaries, which limits their global applicability.
The complexity deepens when models are expected to generalize across languages and cultural domains. Bang et al. (2023) (bang2023enabling, ) and Kaneko et al. (2024) (kaneko2024evaluating, ) underscore the difficulty of maintaining robustness and fairness when supervision lacks cross-linguistic and cross-cultural grounding. These challenges call for inclusive training data that reflects diverse values, usage contexts, and linguistic structures.
To address this, future research should pursue more comprehensive datasets that improve the granularity of safety signals and reduce blind spots in alignment. Such data enables models to make more context-aware judgments and minimize unintentional harm.
7. Conclusions
Through our comprehensive systematic review of LLM challenges in harmful content generation, we have illuminated a rapidly evolving landscape that demands immediate attention and action. Our analysis has uncovered a concerning progression from simple unintentional vulnerabilities to sophisticated intentional exploits, with challenges manifesting across technical, social, and ethical dimensions. While the field has made notable progress in implementing safety measures, the dynamic nature of social media environments and the emergence of advanced attack methods, particularly in multi-modal contexts, continue to present significant challenges. The gap between rapidly advancing LLM capabilities and effective safety measures remains a critical concern, highlighting the pressing need for more resilient and adaptable solutions. As we look to the future of LLM safety, success will depend on our ability to develop proactive, holistic approaches that combine technical innovation with thoughtful policy frameworks, ensuring these powerful language models can serve society while effectively mitigating their potential for harm. This will require sustained collaboration among researchers, industry stakeholders, and policymakers to create and implement multi-stakeholder solutions that can adapt to emerging threats while preserving the beneficial aspects of LLM technology.